# KKL Observer Synthesis for Nonlinear Systems via Physics-Informed Learning

### v1

@@ 7 Simulation Results
In this section, we demonstrate the effectiveness of our proposed methodology for the learning-based design of KKL observers using the following examples:
  • Reverse Duffing oscillator: The state $x\in\mathbb{R}^{2}$ and the dynamics is given by
@@
  • Van der Pol oscillator: The state $x\in\mathbb{R}^{2}$, a parameter $\mu\in\mathbb{R}$, and the dynamics is given by
@@
  • Rössler attractor: The state $x\in\mathbb{R}^{3}$, parameters $a,b,c\in\mathbb{R}$, and the dynamics is given by
    $\begin{array}[]{ccl}\dot{x}_{1}&=&-x_{2}-x_{3}\\ \dot{x}_{2}&=&x_{2}+ax_{2}\\ \dot{x}_{3}&=&b+x_{3}(x_{1}-c)\\ y&=&x_{2}.\end{array}$ (41)
  • Lorenz attractor: The state $x\in\mathbb{R}^{3}$, parameters $p,q,r\in\mathbb{R}$, and the dynamics is given by
    $\begin{array}[]{ccl}\dot{x}_{1}&=&p(x_{2}-x_{1})\\ \dot{x}_{2}&=&x_{1}(q-x_{3})-x_{2}\\ \dot{x}_{3}&=&x_{1}x_{2}-rx_{3}\\ y&=&x_{2}.\end{array}$ (42)
@@
Figure 2: Output prediction with learned KKL observer in the presence of model uncertainty and noisy measurements. The solid black line shows the true state trajectory, while the dashed purple line is the state estimate obtained from the learned KKL observer.
@@ 7.1 Experimental Setup
We generate synthetic datasets and train models for each of the systems (39)–(42) according to our proposed methodology in Section 4. For all systems, we construct the dataset $(X,Z)$ by sampling initial conditions over $[-1,1]^{n_{x}}$ using Latin hypercube sampling, with $n_{x}$ denoting the system dimension, which is 2 and 3 for (39)–(40) and (41)–(42), respectively. We sampled 100 initial conditions for (39), (40), and (42), while we used 200 samples for (41). From each sampled initial condition, we generate system trajectories using Runge-Kutta 4 over a time interval of $[0,50]$.
@@
System Layers Layer Size Learning Rate Epochs
Duffing 3 150 $1e^{-3}$ 15
VdP 2 350 $1e^{-3}$ 15
Rössler 3 250 $1e^{-3}$ 15
Lorenz 2 350 $1e^{-3}$ 15
@@
Table 1: Hyper-parameters used for training. All values are used for both the forward map $\hat{\mathcal{T}}_{\theta}$ and the inverse map $\hat{\mathcal{T}}_{\eta}$.
7.2 Results
@@
Simulation results for each system are illustrated in Fig. 2. Each trajectory is generated from a random initial condition, different from the ones in the dataset $X$, and each component of the model uncertainty and measurement noise is sampled from $\mathcal{N}(0,0.1)$ for the reverse Duffing oscillator, van der Pol oscillator and Rössler attractor, whereas for the Lorenz attractor, it is sampled from $\mathcal{N}(0,2)$. The simulations demonstrate the capability of our method to estimate the unmeasured states with remarkable accuracy even in the presence of model uncertainties and measurement noise.
8 Application to Fault Detection and Isolation
In this section, we demonstrate the effectiveness of the learned KKL observer in detecting and isolating sensor faults in nonlinear systems. We consider two types of sensor faults: failure, where a sensor stops transmitting measurement data, and degradation, where the measurement noise level of a sensor increases. For these faults, the output equation (20b) can be rewritten as
$y(t)=\phi(t)\left[h(x(t))+v(t)+\zeta(t)\right]$ (43)
where
$\phi(t)=\text{diag}(\phi_{1}(t),\dots,\phi_{n_{y}}(t)),\quad\phi_{i}(t)\in\{0,1\}$
models the failure of sensor $i$ when $\phi_{i}(t)=0$, and
$\zeta(t)=[\begin{array}[]{ccc}\zeta_{1}(t)&\dots&\zeta_{n_{y}}(t)\end{array}]^{\top}\in\mathbb{R}^{n_{y}}$
models the degradation of sensor $i$, affecting its measurement accuracy, when $\zeta(t)\neq 0$. Note that a system with fault-free sensors at time $t$ will have $\phi(t)=I_{n_{y}}$ and $\zeta(t)=0_{n_{y}}$. Any other value of these signals would represent a degradation when $\zeta_{i}(t)\neq 0$, for $t\geq t_{i}^{\text{fault}}$, or a failure when $\phi_{i}(t)=0$, for $t\geq t_{i}^{\text{fault}}$, where $i\in[n_{y}]$ and $t_{i}^{\text{fault}}\in\mathbb{R}_{>0}$ is the time at which a fault occurs in sensor $i$.
Given measured output (43) from fault-prone sensors, the output predicted by the learned KKL observer (19) is given by
$\hat{y}(t)=h(\hat{x}(t)).$ (44)
Observer-based fault detection and isolation (FDI) of nonlinear systems relies on the design of observers that must exhibit accurate state estimation and output prediction performance. Using the learned KKL observer (19) for FDI, we address the following problems in this section:
  1. (P1)
    Detection: How do we detect whenever a fault occurs in one or more sensors?
  2. (P2)
    Isolation: How to identify which sensor is faulty after a fault is detected?
To solve these problems, we define residuals to compare the measured output of the system with the predicted output of the learned KKL observer (19). The residual $r_{i}$ corresponding to the $i$-th sensor is defined as
$\begin{array}[]{ccl}r_{i}(t)&\coloneqq&|y_{i}(t)-\hat{y}_{i}(t)|\\ &=&|\phi_{i}(t)[h_{i}(x(t))+v_{i}(t)+\zeta_{i}(t)]-h_{i}(\hat{x}(t))|\end{array}$ (45)
where $y_{i}(t)$ and $\hat{y}_{i}(t)$ are the $i$-th components of $y(t)$ in (43) and $\hat{y}(t)$ in (44), respectively.
Furthermore, consider an increasing sequence of discrete time samples $t_{0}<t_{1}<t_{2}<\dots$ with $\delta_{k}=t_{k}-t_{k-1}$ for $k\in\mathbb{N}$. Then, we define a differentiated residual $\tilde{r}_{i}$ as the absolute value of the numerical differentiation of the residual $r_{i}$, i.e.,
$\tilde{r}_{i}(t_{k})\coloneqq\left|\frac{r_{i}(t_{k})-r_{i}(t_{k-1})}{\delta_{k}}\right|.$ (46)
Let $\tau_{i}\in\mathbb{R}_{>0}$ denote the threshold for residual $r_{i}$ such that, in the steady state, $\limsup_{t\to\infty}r_{i}(t)\leq\tau_{i}$ when there are no faults. Similarly, let $\tilde{r}_{\Delta}\in\mathbb{R}_{>0}$ denote the threshold for all differentiated residuals such that $\limsup_{t\to\infty}\tilde{r}_{i}(t)\leq\tilde{r}_{\Delta}$ when sensor $i$ is not faulty.
8.1 Fault Detection
We derive upper bounds on the residuals $r_{i}(t)$.
Proposition 19.
Suppose Assumptions 10 and 15 hold and the matrix $A$ is diagonalizable with eigenvalue decomposition $A=V\Lambda V^{-1}$. Then, in a fault-less case,
$r_{i}(t)\leq\overline{v}+\ell_{h_{i}}\bigg{[}\epsilon^{*}+ae^{\lambda_{\min}(A)t}+b(t)(\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v})\bigg{]}$ (47)
where $\overline{w},\overline{v},\psi$ are from Assumption 10, $\ell_{h_{i}}$ is the Lipschitz constant of $h_{i}(x)$ for $x\in\mathcal{X}$, $\epsilon^{*}$ is defined in (30), and $a$ and $b(t)$ are given in (34).
{pf}
In a fault-less case, $\phi(t)=I_{n_{y}}$ and $\zeta(t)=0_{n_{y}}$. Thus, from (45),
$r_{i}(t)$ $=|h_{i}(x(t))-h_{i}(\hat{x}(t))+v_{i}|$
$\leq\ell_{h_{i}}\|x(t)-\hat{x}(t)\|+\overline{v}.$
Then, using (33), we obtain (47).
Remark 20.
In the right-hand side of (47), the approximation error bound $\epsilon^{*}$ is an unknown quantity. However, it can be numerically estimated as $\hat{\epsilon}^{*}$ by generating test data and evaluating the maximum error $\|x-\hat{\mathcal{T}}_{\hat{\eta}^{\star}}^{*}(\hat{\mathcal{T}}_{\hat{\theta}^{\star}}(x))\|$ where $x$ are points in the test dataset. $\diamond$
We use Proposition 19 for computing the threshold $\tau_{i}(t)$ on the residuals since
$\limsup_{t\to\infty}r_{i}(t)$
$\qquad\leq\overline{v}+\ell_{h_{i}}\bigg{[}\hat{\epsilon}^{*}+\ell_{\eta}\frac{\text{cond}(V)}{|\lambda_{\min}(A)|}\|B\|(\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v})\bigg{]}$
where $\hat{\epsilon}^{*}$ is an estimate of $\epsilon^{*}$. Therefore, on the test dataset, we can choose the threshold as
$\tau_{i}=\overline{v}+\ell_{h_{i}}\bigg{[}\hat{\epsilon}^{*}+\ell_{\eta}\frac{\text{cond}(V)}{|\lambda_{\min}(A)|}\|B\|(\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v})\bigg{]}.$ (48)
In a steady state when $t$ becomes large, it holds that $r_{i}(t)\leq\tau_{i}$ when there are no sensor faults. When the residual $r_{i}(t)$ surpasses $\tau_{i}$, it signifies that the measured output is significantly different from the predicted output. Since $\tau_{i}$ is the threshold in a fault-free case, sensor faults can be detected by measuring the residuals $r_{i}(t)$ and raising an alert whenever $r_{i}(t)>\tau_{i}$ for any $i$ and any $t\in\mathbb{R}_{\geq 0}$.
8.2 Fault Isolation
A sensor fault is detected whenever the residual $r_{i}(t)$ surpasses the threshold $\tau_{i}$. However, $r_{i}(t)$ exceeding $\tau_{i}$ and having the highest value among other residuals do not mean that sensor $i$ is faulty. This is because a fault in sensor $j$ distorts the output $y_{j}(t)$, which is filtered through the observer dynamics and then transformed by $(h_{j}\circ\hat{\mathcal{T}}_{\eta}^{*})(\cdot)$. Since the neural network $\hat{\mathcal{T}}_{\eta}^{*}$ is not necessarily a diagonal function, the fault in sensor $j$ can induce a large residual $r_{i}(t)$, $i\neq j$, even when sensor $i$ is not faulty. This phenomenon can also be observed in Fig. 3(c) and 4(c), which are discussed below. Due to interdependencies caused by the non-diagonal structure of $\hat{\mathcal{T}}_{\eta}^{*}$, the transients after the occurrence of a fault may persist above the threshold, which causes difficulty in precisely isolating sensor faults merely from the residuals $r_{i}(t)$.
Based on our empirical observations, differentiated residuals $\tilde{r}_{i}(t_{k})$ are better suited for fault isolation. A spike is generated in the differentiated residual $\tilde{r}_{i}(t_{k})$ whenever an abrupt fault ($\phi_{i}(t)=0$ or $\zeta_{i}(t)\neq 0$) is introduced in sensor $i$ at time $t_{k}$. The explanation for this spike is straightforward because, in the differentiated residual, the fault in the output also gets numerically differentiated. Therefore, an abrupt fault results in a spike in the differentiated residual. However, when the fault signal $\zeta_{i}(t)$ is smooth and of low magnitude, it is difficult to isolate it. Nevertheless, in many cases, faults are irregular, non-smooth signals.
We generate test datasets, where we denote state trajectories $x^{j}(t_{k})\coloneqq x(t_{k};x_{0}^{j},0)$ that are initialized at $q$ initial points $x(0)=x_{0}^{j}$, for $j=1,\dots,q$, and the corresponding observer estimates as $\hat{x}^{j}(t_{k})\coloneqq\hat{x}(t_{k};\hat{\mathcal{T}}_{\eta}^{*}(z_{0}^{j}))$. We simulate the KKL observer and compute the residuals $r_{1}^{j}(t_{k}),\dots,r_{n_{y}}^{j}(t_{k})$, for $k=0,\dots,T$, where $T\in\mathbb{N}$ and
$r_{i}^{j}(t_{k})=|y_{i}(t_{k};x^{j})-\hat{y}_{i}(t_{k};\hat{x}^{j})|.$
Here, $t_{T}>0$ is chosen large enough to allow the observer to converge to a neighborhood of the original state. Then, the corresponding differentiated residuals $\tilde{r}_{1}^{j}(t_{k}),\dots,\tilde{r}_{n_{y}}^{j}(t_{k})$, for $k=1,\dots,T$, are obtained using (46). Let $t_{c}$, for $c<T$, denote the minimum time at which the observer is in the steady state. Then an empirical threshold for differentiated residuals is computed on the test dataset as follows:
$r_{\Delta}=\max_{\begin{subarray}{c}k\in\{c,\dots,T\}\\ j\in[q]\\ i\in[n_{y}]\end{subarray}}\tilde{r}_{i}^{j}(t_{k}).$ (49)
Given that the learned KKL observer estimates the state of the system accurately, $\tilde{r}_{i}(t)\leq r_{\Delta}$ remains steady when $t_{k}$ becomes sufficiently large. Therefore, whenever the differentiated residual $\tilde{r}_{i}(t)$ surpasses the empirical threshold $r_{\Delta}$ given by (49), an alert can be raised that sensor $i$ is faulty. We validate this observation via simulations in the next section.
8.3 Network of Kuramoto Oscillators
We consider a network of Kuramoto oscillators and show that FDI of a highly nonlinear system can be achieved using the learned KKL observer (19), even in the presence of model uncertainties and measurement noise. Both sensor failure and degradation faults are demonstrated.
The network of Kuramoto oscillators describes synchronization phenomena in interconnected systems and has applications in electrical power networks, multi-agent coordination, and distributed software systems [18]. The dynamics of a network of Kuramoto oscillators with $n$ nodes is given as
$\dot{\theta_{i}}(t)=\omega_{i}+\sum_{j=1}^{n}a_{ij}\sin(\theta_{i}(t)-\theta_{j}(t))$ (50)
where $\theta_{i}(t)\in\mathbb{S}^{1}$ is the phase angle of node $i=1,\dots,n$, $\omega_{i}\in\mathbb{R}$ is the natural frequency of node $i$, and $a_{ij}\geq 0$ denotes the coupling between node $i$ and $j$. In the literature, the state trajectories of (50) are often represented graphically as $\sin(\theta_{i})$ to better illustrate their synchronization. We follow the same convention in our simulations.
8.4 Experimental Setup
We consider a network of $n=10$ nodes with randomly generated natural frequencies $\omega_{i}$ and couplings $a_{ij}$. The measurements are chosen as $y=\begin{bmatrix}\theta_{1}&\theta_{2}&\theta_{3}&\theta_{4}&\theta_{5}\end{bmatrix}^{\top}$. A set of 100 initial conditions is generated in $[-2,2]^{10}$ using Latin hypercube sampling. We choose Runge-Kutta as our numerical ODE solver to simulate (50) and (7) over a time interval $[0,30]$, partitioned into $4000$ sample points for each trajectory. The neural network $\hat{\mathcal{T}}^{*}_{\eta}$ is chosen as a fully connected feed-forward network, consisting of 3 hidden layers of 250 neurons with ReLU activation functions. Model training is facilitated by data standardization and learning rate scheduling. Following [9], the parameters of KKL observer (19) are chosen as
$A=\Lambda\otimes I_{n_{y}},\qquad B=\Gamma\otimes I_{n_{y}}$
where $\Lambda\in\mathbb{R}^{(2n_{x}+1)\times(2n_{x}+1)}$ is a diagonal matrix with diagonal elements linearly distributed in $[-15,-21]$, $\Gamma\in\mathbb{R}^{2n_{x}+1}$ is a column vector of ones, and $I_{n_{y}}$ is the identity matrix of size $n_{y}\times n_{y}$. Here, $n_{x}$ and $n_{y}$ are 10 and 5, respectively, and $n_{z}=n_{y}(2n_{x}+1)=105$.
(a) Output prediction
(b) Fault signal
(c) Residuals
(d) Differentiated residuals
Figure 3: Detection and isolation of sensor $4$’s failure at time $10$.
(a) Output prediction
(b) Fault signal
(c) Residuals
(d) Differentiated residuals
Figure 4: Detection and isolation of sensor 1 and 5’s degradation caused by bias at times 5 and 15, respectively.
(a) Output prediction
(b) Fault signal
(c) Residuals
(d) Differentiated residuals
Figure 5: Detection and isolation of sensor 3’s degradation caused by increasing noise over time.
8.5 Results
Figures 35 show the results of FDI for a network of Kuramoto oscillators using the learned KKL observer. The theoretical thresholds $\tau_{i}$ of the residuals $r_{i}(t)$ computed by (48) are shown in the figures in the third column, while the empirical threshold $r_{\Delta}=4.74$ is computed using (49) according to the method described in Section 8.2. Going from left to right, the figures show (a) the predicted output $\hat{y}_{i}$ obtained by the learned KKL observer and the true output $y_{i}$ measured by the faulty sensors, (b) the active fault-inducing signals $\phi_{i}(t)$ or $\zeta_{i}(t)$, (c) the generated residuals $r_{i}$ and their respective thresholds $\tau_{i}$ and (d) the differentiated residuals $\tilde{r}_{i}$ and their common threshold $\tilde{r}_{\Delta}$.
Fig. 3 shows that our method is capable of detecting and isolating sensor failures, which is demonstrated by inducing sensor 44{4}4’s failure with $\phi_{4}(t)=0$ for $t\geq 10$. Before the fault, Fig. 3(a) shows the observer successfully tracking the noisy output of the system after a short transient period induced by the initialization $\hat{z}(0)=0$. The introduction of sensor 4’s failure at time $t=10$ (Fig. 3(b)) causes the predicted output to diverge from the actual output, producing large residuals that are observed in Fig. 3(c). Although sensor 4 failed, we notice in Fig. 3(c) that the residuals corresponding to sensor 3 and 5 crossed their respective thresholds, whereas the residual corresponding to sensor 4 remained under the threshold. This indicates that monitoring the residuals can allow fault detection but not isolation. For fault isolation, we monitor differentiated residuals in Fig. 3(d) and observe the spike in sensor 4’s differentiated residual at $t=10$, accurately indicating its failure.
Fig. 4 and 5 demonstrate the detection and isolation of sensor degradation. In Fig. 4(b), sensors 11{1}1 and 55{5}5 are introduced with a bias $\zeta_{1}(t)=1$, at $t\geq 5$, and $\zeta_{5}(t)=1$ at $t\geq 15$. Fig. 4(a) shows output prediction. Each fault is detected at the moment of occurrence from the residuals in Fig. 4(c) and is isolated using the differentiated residuals in Fig. 4(d). Lastly, Fig. 5(b) shows the fault signal on sensor $3$, which is a gradually increasing white noise. Fig. 5(a) shows output prediction and Fig. 5(c) and Fig. 5(d) illustrate fault detection and isolation, respectively.

### v2

@@ 8 Simulation Results
In this section, we demonstrate the effectiveness of our proposed methodology [^7] Our code: github.com/Mudhdhoo/PINN_KKL for the learning-based design of KKL observers using the following examples:
  • Reverse Duffing oscillator with state $x\in\mathbb{R}^{2}$:
@@
  • Van der Pol oscillator with state $x\in\mathbb{R}^{2}$ and parameter $\mu=3$:
@@
  • Rössler attractor with state $x\in\mathbb{R}^{3}$ and parameters $a=b=0.2$ and $c=5.7$:
    $\begin{array}[]{cclccl}\dot{x}_{1}&=&-x_{2}-x_{3},&\dot{x}_{2}&=&x_{1}+ax_{2}\\ \dot{x}_{3}&=&b+x_{3}(x_{1}-c),&y&=&x_{2}.\end{array}$ (39)
  • Lorenz attractor with state $x\in\mathbb{R}^{3}$ and parameters $p=28$, $q=10$, and $r=8/3$:
    $\begin{array}[]{cclccl}\dot{x}_{1}&=&p(x_{2}-x_{1}),&\dot{x}_{2}&=&x_{1}(q-x_{3})-x_{2}\\ \dot{x}_{3}&=&x_{1}x_{2}-rx_{3},&y&=&x_{2}.\end{array}$ (40)
@@
Figure 2: State estimation with learned KKL observer under noisy measurements. The solid black line shows the true state trajectory, while the dashed magenta line is the state estimate obtained from the learned KKL observer.
@@ 8.1 Experimental Setup
We generate synthetic datasets $S_{1}$ and $S_{2}$ according to the procedure described in Sections 5.2 and 5.3, and train neural networks for each of the systems (37)–(40) according to our proposed methodology in Section 5. For all systems, we construct the dataset $S_{1}$ by uniformly sampling initial conditions over $1,1]^{n_{x}}$, with $n_{x}$ denoting the system dimension, which is 2 and 3 for (37)–(38) and (39)–(40), respectively. We sample 100 initial conditions for (37)–(38), while we used 200 samples for (39)–(40). From each sampled initial condition, we generate system trajectories over a time interval of $[0,50]$ with sampling time $0.1$. The truncation sample $k_{*}$ is set to $0.1\tau$ (10% of the $z$-trajectory length), which is sufficient to ensure that the effects of all initial conditions vanish by $\varepsilon=10^{-4}$. The observer matrices $A=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n_{z}})$, where $\lambda_{i}\in[-2,-0.5]$ are distinct, and $B=\mathbf{1}_{n_{z}}$, where $\mathbf{1}_{n_{z}}$ is a vector of ones.
@@
Table 1: Hyperparameters used for training. All values are used for both the forward map $\hat{\mathcal{T}}_{\theta}$ and the inverse map $\hat{\mathcal{T}}_{\eta}$.
System Hidden Layers Layer Size Learning Rate Physics weight $\nu$ Epochs
Duffing 3 150 $\times 10^{-3} 1 1$5
VdP 2 350 $\times 10^{-3} 1 1$5
Rössler 3 250 $\times 10^{-3} 1 1$5
Lorenz 2 350 $\times 10^{-3} 1 1$5
8.2 Estimation Results under Measurement Noise
@@
Simulation results for each system are illustrated in Fig. 2. Each trajectory is generated from random initial condition, different from those in the dataset $S_{1}$ but sampled within the same region. Each component of the measurement noise is sampled from $\mathcal{N}(0,0.1)$ for the reverse Duffing oscillator, van der Pol oscillator, and Rössler attractor. In contrast, for the Lorenz attractor, it is sampled from $\mathcal{N}(0,2)$. In all cases, the observer’s initial condition $\hat{z}(0)=0$. The simulations demonstrate the capability of our method to estimate the state with remarkable accuracy even in the presence of measurement noise.
8.3 Comparison with the State-of-the-Art
We compare the state estimation performance of our sequential learning PINN method (referred to as SPINN) with three established learning-based approaches:
The benchmark system is the reverse Duffing oscillator (37). We use the same neural network architecture and hyperparameters (described in Section 8.1 and Table 1) for all methods to ensure a fair comparison and focus solely on the differences in the learning method. We also assess the training costs of each method, which is a critical factor for scalability and rapid prototyping.
In this comparison, we evaluate the generalization capabilities of the learned observers produced by these methods. All methods are trained on data generated in the set $1,1]^{n_{x}}$. Generalization performance is evaluated by testing over 100 independent trials with initial states of system 37 sampled from a larger set $3,3]^{n_{x}}$, which the neural networks did not see during training. Recall that $x(t)$ is the true state and $\hat{x}(t)$ is the estimated state. We use two standard metrics to quantify the estimation accuracy across the $N_{\mathrm{traj}}=100$ independent test trajectories with the number of time samples $N_{\mathrm{steps}}$:
  • Root Mean Square Error
    $\text{RMSE}=\sqrt{\frac{1}{N_{\mathrm{traj}}N_{\mathrm{steps}}}\sum_{j=1}^{N_{\mathrm{traj}}}\sum_{t=1}^{N_{\mathrm{steps}}}\|x_{j}(t)-\hat{x}_{j}(t)\|^{2}}$
    measures the absolute magnitude of the average error.
  • Symmetric Mean Absolute Percentage Error
    $\text{SMAPE}=\frac{1}{N_{\mathrm{traj}}N_{\mathrm{steps}}}\!\sum_{j=1}^{N_{\mathrm{traj}}}\!\sum_{t=1}^{N_{\mathrm{steps}}}\!\frac{2\|x_{j}(t)-\hat{x}_{j}(t)\|}{\|x_{j}(t)\|+\|\hat{x}_{j}(t)\|}\times 100\%$
    measures the relative estimation error for comparing accuracy across different scales.
Fig. 3 illustrates a single instance of the estimation performance when the system is initialized outside the training domain. It visually confirms that our observer estimates the state with the best accuracy even when the system is initialized from $3,3]^{n_{x}}$, which is larger than the training domain. On the other hand, Fig. 4 presents the quantitative error metrics for the 100 test trials. The results show that the generalization performance of our method (SPINN) is significantly better than all other methods in this out-of-domain generalization test.
Figure 3: Estimation performance of KKL observers learned from different methods when the system is initialized outside the training domain.
Figure 4: Comparison of quantitative error metrics averaged over the 100 trials.
The observed performance hierarchy yields key insights into the value of physics-informed learning. Our method achieves the best generalization because it (a) explicitly incorporates the PDE constraint in (16), which acts as a physically consistent inductive bias avoiding overfitting on training data, and (b) capitalizes on the sequential decoupling, which improves learning by avoiding the conflicting gradients inherent in joint $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ training. The PDE (6) is incorporated into neural network training via automatic differentiation and the known system model $f(x)$, thereby reducing the size of the hypothesis space and reliably facilitating extrapolation into unseen state regions. This observation validates the theoretical arguments of Section 6.
The unsupervised AE [15] and DMF [42] methods also exhibit reasonable generalization performance. NODE [36] performs relatively worse in this test. This is because it handles unknown dynamics simultaneously with the unknown state. Since the system model is learned from limited training data in the domain $1,1]^{n_{x}}$ without any physics knowledge guiding the training in the hypothesis space, the learned model fails to generalize when tested on the unseen domain $3,3]^{n_{x}}$. This model identification error, in turn, contaminates the observer’s performance. Note that it is an expected limitation of simultaneous model identification and state estimation in a limited-data regime. On the other hand, the DMF method [42], which is also model-free, benefits from a streamlined learning objective that directly targets the KKL injectivity property, without inheriting the costly burden of learning the system dynamics from data.
Table 2: Computational times for training.
Method Training Time (s) Relative Cost Computational   Bottleneck
SPINN (our) 18.29 1x Sequential procedure and computation of PDE Jacobian $\nabla_{x}\hat{\mathcal{T}}$
AE 7.68 0.42x Simple regression
NODE 219.97 12.03x Numerical ODE integration per forward pass
DMF 7.69 0.42x Simple regression
All models were trained on a MacBook Pro equipped with the Apple M3 Max, featuring a 16-core CPU (12 performance cores and 4 efficiency cores) and a 40-core GPU. The system utilizes 128 GB of unified memory with a memory bandwidth of 400 GB/s. The total training times for the 15 epochs (or equivalent iterations) are reported in Table 2. The unsupervised AE [15] and deep model-free [42] methods are the most efficient ($\sim$7.7s). Their objective functions rely solely on algebraic reconstruction losses, allowing for extremely fast backpropagation. Our method (SPINN) required approximately $2.4\times$ ($\sim$18.3s) the training time of these methods. This overhead is the direct cost of the physics-informed constraint: evaluating the PDE residual requires computing the Jacobian $\frac{\partial\hat{\mathcal{T}}}{\partial x}$ via automatic differentiation at every step. However, this cost is constant and leads to the significant generalization gains observed in Fig. 4. The neural ODE method [36] is orders of magnitude slower ($\sim$220s). This is because every forward pass of the network requires a numerical ODE solver to integrate the learned neural ODE dynamics over time. Backpropagating through this solver (or using the adjoint method) is computationally expensive, making the NODE method less suitable for scenarios that require rapid retraining or real-time adaptation.